Consider a plane wave incident on a convex lens of diameter 5cm and of focal length 10cm. If the wavelength of the incident light is 6000°A, calculate the radius of the first dark ring on the focal plane of the lens. Repeat the calculation for a lens of same focal length but diameter 15cm. Interpret the result physically.
1
Expert's answer
2019-10-07T11:04:37-0400

When a circular aperture of radius "a" placed in front of a convex lens of focal length "f" and illuminated by a parallel beam of light of wavelength "\lambda" then the radius of the first dark ring is given as follows:

"R=\frac{3.832\lambda f}{2\pi a}"

then, for "a=2.5\, cm"

"R=\frac{3.832\cdot 6\cdot10^{-7}\cdot 0.1}{2\cdot\pi \cdot0.025} = 1.46\cdot 10^{-7} m"

and for "a=7.5\, cm"

"R=\frac{3.832\cdot 6\cdot10^{-7}\cdot 0.1}{2\cdot\pi \cdot0.075} = 4.88\cdot 10^{-8} m"